# Probing Visual Language Priors in VLMs

### v1

@@ Abstract
Despite recent advances in Vision-Language Models (VLMs), many still over-rely on visual language priors existed in their training data rather than true visual reasoning. To examine the situation, we introduce ViLP, a visual question answering (VQA) benchmark that pairs each question with three potential answers and three corresponding images: one image whose answer can be inferred from text alone, and two images that demand visual reasoning. By leveraging image generative models, we ensure significant variation in texture, shape, conceptual combinations, hallucinated elements, and proverb-based contexts, making our benchmark images distinctly out-of-distribution. While humans achieve near-perfect accuracy, modern VLMs falter; for instance, GPT-4 achieves only 66.17% on ViLP. To alleviate this, we propose a self-improving framework in which models generate new VQA pairs and images, then apply pixel-level and semantic corruptions to form “good-bad” image pairs for self-training. Our training objectives compel VLMs to focus more on actual visual inputs and have demonstrated the effectiveness in enhancing the performance of open-source VLMs, including LLaVA-v1.5 and Cambrian.
@@ 1 Introduction
Vision-Language Models (VLMs) have advanced text-image interaction, serving as a cornerstone in bridging the gap between visual and textual data (Achiam et al., 2023; Team et al., 2023). However, a persistent challenge for learning-based models, such as ResNets and CLIPs, is their reliance on strong priors overly learned in the training data, sometimes overlooking visual cues for answering questions (Agrawal et al., 2016; Prabhu et al., 2023). For example, when shown an image of a hexagonal moon (Figure 1), a model might incorrectly assert that “the moon is round" due to strong language priors. Similarly, these models may adhere to learned visual priors (Thrush et al., 2022; Sterz et al., 2024), making it difficult for them to comprehend out-of-distribution visual cues, such as a zebra with atypical spot patterns (Figure 1), where humans would easily discern the difference. This raises an important question: do today’s VLMs still over-rely on learned visual language priors?
@@
To investigate this, we probe the visual language priors of VLMs by constructing Question-Image-Answer (QIA) triplets that deliberately deviate from the training data distribution. Unlike existing benchmarks that typically rely on internet-sourced images, which inadvertently favor the visual language priors embedded in training data, we utilize modern image generation models such as DALL·E-3 (Ramesh et al., 2021) and Flux to synthesize out-of-distribution images, supporting uncommon answers. To further challenge the priors of VLMs, we amplify the language priors in questions by introducing distractor facts, making the questions more suggestive.
@@
Our benchmark, ViLP, contains 300 carefully designed questions, each paired with three distinct answers: a Prior Answer and two Test Answers, resulting in a total of 900 QIA triplets. Each question is structured to present a distractor fact followed by a question. The distractor fact directly leads to the Prior Answer. In contrast, the two Test Answers are crafted to challenge the priors by requiring both textual and visual cues for accurate reasoning. While human participants can differentiate between all three answers, current VLMs face considerable difficulty, as evidenced by a significant performance drop in our benchmarks, with GPT-4o (OpenAI, 2024) scoring only 66.17%.
@@
Inspired by ViLP  results, we propose a self-improving approach to encourage VLM reasoning more relying on visual inputs. Our method involves self-generating VQAs using image generation and editing models (Podell et al., 2023; Ren et al., 2024; Brooks et al., 2023a) and introducing controlled corruptions to create “good" and “bad" image pairs for DPO-like training (Rafailov et al., 2024). We theoretically prove that our proposed objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022), and experiments on open-sourced VLMs, including LLaVA-v1.5 (Liu et al., 2024) and Cambrian (Tong et al., 2024), demonstrate its effectiveness.
@@
Figure 1: Sample data from ViLP. For the same question, ViLP  provides three distinct images and corresponding answers (upper-left corner). All questions follow a consistent structure, combining a distractor fact with a question. The Prior Answer (first column) can be directly inferred from the question, while Test Answers (second &\&& third column) rely on visual cues. All of our answers are designed to be single words, allowing for evaluation without the need for LLMs. We have developed a robust synonym and plural detection pipeline to ensure that open-ended responses do not interfere with the evaluation process. Please refer to Appendix A for more data samples from ViLP.
@@ 2 Related Work
VQA Dataset: There have been long-term efforts to establish Visual Question Answering (VQA) datasets from various perspectives, including general visual question answering  (Agrawal et al., 2015; Gurari et al., 2018; Fu et al., 2023; Liu et al., 2023e; Li et al., 2023a; Yu et al., 2023b; Liu et al., 2024), reading text or charts (Singh et al., 2019a; Mathew et al., 2020; 2021; Masry et al., 2022), complicated reasoning (Lu et al., 2022; 2023), probing compositions (Hudson & Manning, 2019; Ma et al., 2022; Thrush et al., 2022; Hsieh et al., 2023; Li et al., 2024a), studying hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Guan et al., 2023), common sense reasoning (Bitton-Guetta et al., 2023b; a), and more (Majumdar et al., 2024; Sterz et al., 2024). This paper proposed a benchmark to evaluate the visual reasoning abilities of VLMs when the questions, answers and images defy common visual language priors. Our approach follows the balanced dataset design in Goyal et al. (2017), where each question is associated with three answers: the first aligns with language priors, while the latter two deviate from them, requiring reliance on visual cues for correct answering. By leveraging state-of-the-art image generation models, our benchmark creates images that defy priors much more effectively than previous works using internet images, such as Goyal et al. (2017); Tong et al. (2024). Moreover, unlike the “trick” category in Sterz et al. (2024), our process first generates question-answer pairs and then synthesizes images that meet the specified conditions, resulting in more challenging visual reasoning tasks. More comparisons with existing datasets are included in Appendix D.2.
@@
Vision Language Models and Language Priors: Multimodal reasoning is essential for machine intelligence, with vision-language models (VLMs) integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b; a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c; a; Zhao et al., 2023; Chen et al., 2023; Zhu et al., 2024; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024). Building on the success of language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a; b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent advancements leverage vision-language paired datasets (Liu et al., 2024; Tong et al., 2024) to fine-tune intermediate connectors between LLMs and visual encoders (Liu et al., 2024). However, these paired datasets are significantly smaller than the large text corpora used for LLM pre-training (OpenAI, 2023b; Soldaini et al., 2024), and many approaches do not update the visual encoder and LLM parameters during fine-tuning, which can lead to persistent language biases. Specifically, the outputs of VLMs become dominated by the language priors learned from the pre-training corpus, reducing the importance of visual inputs (Thrush et al., 2022; Sterz et al., 2024). This issue becomes even more critical when VLMs are presented with deliberately generated images that challenge these biases, as demonstrated in our work. Previous methods and datasets (Goyal et al., 2016; Agrawal et al., 2017; Dancette et al., 2021; Wu et al., 2022; Ramakrishnan et al., 2018; Gouthaman & Mittal, 2020) have attempted to assess these issues by studying visual backbones and curating datasets using simulators (Johnson et al., 2016) or internet images (Zhang et al., 2016). This paper examines these challenges by creating a novel VQA benchmark, featuring carefully designed questions, fact-based distractors, and rare-distribution answers. Furthermore, we utilize advanced image generation techniques to produce realistic images that subvert common language biases as shown in Figure 1.
Self-Rewarding VLM: Self-rewarding LLM (Yuan et al., 2024) has demonstrated the potential of using LLMs to generate and optimize data through directed reference optimization (Rafailov et al., 2024), leading to self-improvements. This line of research has been extended to VLMs, where new answers are generated and rated for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024c; Wang et al., 2024b; a). Our findings resonate with this body of work on self-rewarding VLMs; however, we present two key distinctions: (1) our proposed image-DPO method generates multiple images for a single question-image pair, rather than generating multiple answers and associated ratings; (2) unlike previous methods that focused solely on existing images, both our proposed image-DPO and text-DPO pipelines produce a diverse array of new images with pre-trained image generative models, including SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), and InstructPix2Pix (Brooks et al., 2023a). Specifically, as illustrated in Figure 6, our proposed text-DPO is similar to the self-rewarding VLMs mentioned earlier but incorporates the generation of new images using pre-trained image generation models. In contrast, our image-DPO deliberately corrupts images to create multiple degraded versions, which are used as rejected data during DPO training. This is intended to compel the VLM to rely more on the visual input for accurate generation.
@@ 3.1 Design Principles
"What’s the tall animal with a long neck?" Even without any visual cues, humans readily infer that the answer is "giraffe", leveraging language context and commonsense knowledge. But is it necessarily a giraffe? As shown in the bottom-right of Figure 1, there may be an elephant or a camel with a very long neck and much taller than giraffes. This example highlights a potential issue with current large Vision-Language Models (VLMs). As VLMs are built upon Large Language Models (LLMs) trained on web-scale corpora and are fine-tuned with comparatively much smaller image-text datasets, they may develop a strong bias toward certain visual language priors (especially language) presented in training data and overlook real visual input (Thrush et al., 2022; Sterz et al., 2024). Regarding visual language priors, we focus on strong language priors that confuse VLMs to derive direct answers solely from textual cues. Simultaneously, we aim to create images that are distinctly out-of-distribution to challenge potential visual priors which make the model overlook the critical visual cues such as the unusual spots patterns in the Zebra shown in Figure 1.
To assess the visual reasoning abilities of VLMs against visual language priors, we introduce a new benchmark, ViLP, featuring carefully constructed out-of-distribution question-image-answer (QIA) triplets. These triplets are designed based on the below principles ensuring that VLMs rely on both textual and visual cues, rather than learned priors, to arrive at correct answers.
  • Text-Only Inference: Create questions that, when considered alone, lead to a high-confidence answer based solely on textual information.
  • Visual Influence: When an accompanying image is present, the correct answer differs significantly from the text-only response, sometimes contradicting common sense, ensuring the critical role of visual context in ensuring a high-confidence result.
Mathematically, with $Q$ denoting a question, $I$ denoting an image, $A=\{a_{\text{prior}},\cdots,a_{\text{test}},\cdots\}$ being the set of all possible answers, we use $P(a\mid Q)$ and $P(a\mid Q,I)$ to denote the probability of answer $a$ given the question $Q$ alone, and the probability of answer $a$ given both the question $Q$ and the image $I$, respectively. Here $p$ is a prior model, which could either be human’s cognition prior ($P_{\text{human}}$) or a VLM/LLM’s learned visual language prior ($P_{\theta}$). For constructing our benchmark, we used the following objectives:
Criteria One: Without considering the image, the question $Q$ should lead to a high-confidence answer $a_{\text{prior}}$, where $\delta_{1}$ is a high-confidence threshold. $a_{\text{prior}}$ usually satisfies common knowledge, such as “the moon is round" and “Leonardo da Vinci painted the Mona Lisa" as shown in Figure 1.
@@
Criteria Two: With the inclusion of the image $I$, the correct answer becomes $a_{\text{test}}$, where $\delta_{2}$ is another high-confidence threshold. Also, $a_{\text{test}}$ is significantly different from $a_{\text{prior}}$ to ensure that the image has a substantial impact on the answer, where $D$ is a divergence measure (e.g., Kullback-Leibler divergence), and $\delta_{3}$ is a threshold indicating significant difference. For instance, the image in the second-to-last row and column of Figure 1 clearly shows the moon as a hexagon.
$P(a_{\text{test}}\mid Q,I)\geq\delta_{2}\quad\text{and}\quad D\big{(}P(a_{\text{prior}}\mid Q),P(a_{\text{test}}\mid Q,I)\big{)}\geq\delta_{3}$ (2)
Criteria Three: We design answer $a_{\text{test}}$ to be a rare choice, one that is unlikely to be inferred from the question $Q$ alone. In contrast, answer $a_{\text{prior}}$ should be clearly incorrect when considering both the question $Q$ and the image $I$, where $\delta_{4}$ represents a low-confidence threshold. Typically, $a_{\text{test}}$ will be out of context. For example, in the image from the 3rd row and 2nd column of Figure 1, Vincent van Gogh serves as $a_{\text{test}}$, inferred from the image, while Leonardo becomes contradictory after incorporating the additional image condition.
$P(a_{\text{test}}\mid Q)\leq\delta_{4}\quad\text{and}\quad P(a_{\text{prior}}\mid Q,I)\leq\delta_{4}$ (3)
In constructing ViLP, we use human cognition prior $P_{\text{human}}$ as guidance for its design. We evaluated the behavior of the learned priors of VLMs/LLMs by benchmarking $P_{\theta}$ against the human prior $P_{\text{human}}$. This benchmarking process implicitly measuring the divergence between $P_{\text{human}}$ and $P_{\theta}$.

### v3

@@ Abstract
Despite recent advances in Vision-Language Models (VLMs), they may over-rely on visual language priors existing in their training data rather than true visual reasoning. To investigate this, we introduce ViLP, a benchmark featuring deliberately out-of-distribution images synthesized via image generation models and out-of-distribution Q&A pairs. Each question in ViLP is coupled with three potential answers and three corresponding images: one that can be resolved by text priors alone and two that demand visual reasoning. Although, humans achieve near-perfect accuracy, modern VLMs falter; for instance, GPT-4 achieves only 66.17% on ViLP. To alleviate this, we propose a self-improving framework in which models generate new VQA data, then apply pixel-level and semantic corruptions to form “good-bad" image pairs for self-training. Our training objectives compel VLMs to focus more on the actual visual inputs, and we demonstrate their effectiveness in boosting the performance of open-source VLMs, including LLaVA-v1.5 and Cambrian.
Machine Learning, ICML
[…]
@@ 1 Introduction
Vision-Language Models (VLMs) have advanced text-image interaction, bridging the gap between visual and textual data (Achiam et al., 2023; Team et al., 2023). However, a persistent challenge for learning-based models, such as ResNets and CLIPs, lies in their reliance on learned priors from the training data, sometimes overlooking visual cues when answering questions (Agrawal et al., 2016; Prabhu et al., 2023). For example, when shown a torus-shaped soccer ball (Figure 1), a model might incorrectly identify it as a sphere due to strong language priors. Simultaneously, these models may adhere to learned visual priors (Thrush et al., 2022; Sterz et al., 2024), making it difficult to comprehend out-of-distribution visual cues, such as a zebra with atypical spot patterns (Figure 1), which humans would easily discern. This raises an important question: do today’s VLMs still over-rely on learned visual language priors, especially given that they rely on far fewer image-text pairs compared to the internet-scale text corpora used for pretraining?
@@
To investigate this, we probe the Visual Language Priors of VLMs by constructing Question-Image-Answer (QIA) triplets that deliberately deviate from the training data distribution. Unlike existing benchmarks that typically rely on internet-sourced images (Goyal et al., 2017; Tong et al., 2024), which inadvertently favor the visual language priors embedded in the training data of VLMs, we utilize modern image generation models, including DALL·E-3 (Ramesh et al., 2021) and Flux, to synthesize out-of-distribution images supporting also out-of-distribution answers. Using the image generation models also helps our QIAs exhibit notable variation in texture, shape, conceptual combinations, hallucinated elements, and proverb-based contexts.
@@
Our benchmark, ViLP, contains 300 carefully designed questions, each paired with three distinct answers: a Prior Answer and two Test Answers, resulting in a total of 900 QIA triplets. To further challenge the priors of VLMs, we amplify language priors in questions by introducing distractor facts: each question is structured to present a distractor fact followed by a question. The distractor fact directly leads to the Prior Answer. In contrast, the two Test Answers are crafted to challenge the priors by requiring both textual and visual cues for accurate reasoning. While human participants achieved $\sim$98accuracy easily, current VLMs exhibit considerable difficulty, as evidenced by a significant performance drop on our benchmarks, with GPT-4o (OpenAI, 2024) scoring only 66.17%.
@@
Motivated by the results of ViLP, we propose Image-DPO, a self-improving approach to enhance VLM visual reasoning performance by increasing reliance on visual inputs. Our method employs self-generated VQAs using image generation and editing models (Podell et al., 2023; Ren et al., 2024; Brooks et al., 2023a) and applies controlled corruptions to create “good" and “bad" image pairs for DPO-like training (Rafailov et al., 2024). Experiments with open-source VLMs, including LLaVA-v1.5 (Liu et al., 2024a) and Cambrian (Tong et al., 2024), demonstrate its effectiveness. Moreover, we theoretically show that our objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022). The proposed ViLP dataset, Image-DPO code, and our synthetic data appear in https://huggingface.co/ViLP.
@@
Figure 1: Sample data from ViLP. For the same question, ViLP  provides three distinct images and corresponding answers (upper-left corner). All questions follow a consistent structure, combining a distractor fact with a question. The Prior Answer (first column) can be directly inferred from the question, while Test Answers (second &\&& third column) rely on visual cues. Our answers are designed to be single words, and both the model and human evaluators are tasked with open-domain answering, rather than selecting from predefined options. To support this, we have developed a robust synonym and plural detection pipeline, ensuring that open-ended responses do not hinder the evaluation process. This approach also enables evaluation without relying on LLMs. Please refer to Appendix A.1 for more data samples from ViLP. Furthermore, we include both qualitative and quantitative comparison results with Winoground (Thrush et al., 2022), Whoops!(Bitton-Guetta et al., 2023a), and HallusionBench(Guan et al., 2023) in Appendix A.3.
@@ 2 Related Work
VQA Dataset: Significant efforts have produced VQA datasets from various angles, including general VQA (Agrawal et al., 2015; Gurari et al., 2018; Fu et al., 2023; Liu et al., 2023d; Li et al., 2023a; Yu et al., 2023b; Liu et al., 2024a), reading text or charts (Singh et al., 2019a; Mathew et al., 2020, 2021; Masry et al., 2022), complex reasoning (Lu et al., 2022, 2023), composition probing (Hudson & Manning, 2019; Ma et al., 2022; Thrush et al., 2022; Hsieh et al., 2023; Li et al., 2024a), hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Guan et al., 2023), common-sense reasoning (Bitton-Guetta et al., 2023b, a), and more (Majumdar et al., 2024; Sterz et al., 2024). We propose a benchmark that tests VLMs’ visual reasoning when questions, answers, and images defy common patterns. Following the balanced dataset design of (Goyal et al., 2017), each question is accompanied by three answers: one aligns with language priors, and two deviate, prompting reliance on visual cues. By leveraging state-of-the-art image generation models, our benchmark challenges these priors more effectively than previous datasets built from internet images (Goyal et al., 2017; Tong et al., 2024). Furthermore, unlike the “trick” category in (Sterz et al., 2024), we first generate question-answer pairs before synthesizing images under specified constraints, creating more difficult visual reasoning tasks. Comprehensive comparisons with existing datasets appear in Appendix A.3.
@@
Vision Language Models and Language Priors: Multimodal reasoning is crucial for machine intelligence, with VLMs integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b, a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c, a; Zhao et al., 2023a; Chen et al., 2023; Zhu et al., 2024; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024). Inspired by the success of large language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a, b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent methods leverage relatively small vision-language paired datasets (Liu et al., 2024a; Tong et al., 2024) to fine-tune connectors between LLMs and visual backbones (Liu et al., 2024a). However, these datasets are far smaller than the vast text corpora for LLM pre-training (OpenAI, 2023b; Soldaini et al., 2024), and freezing the visual encoder and LLM parameters often preserves language biases, causing visual inputs to be overshadowed (Thrush et al., 2022; Sterz et al., 2024). This challenge is amplified by deliberately generated images that expose such biases, as shown in our study. Previous works and datasets (Goyal et al., 2016; Agrawal et al., 2017; Dancette et al., 2021; Wu et al., 2022; Ramakrishnan et al., 2018; Gouthaman & Mittal, 2020) addressed these issues with curated simulators (Johnson et al., 2016) or internet imagery (Zhang et al., 2016). In this paper, we present a novel VQA benchmark featuring carefully designed questions, fact-based distractors, rare-distribution answers, and image generation techniques to produce realistic visuals that challenge learned visual language priors (Figure 1).
Self-Rewarding VLM: Self-rewarding LLM (Yuan et al., 2024) has shown that LLMs can generate and improving themselves in the process via Directed Peference Optimization (DPO) (Rafailov et al., 2024). This approach extends to VLMs by generating new answers for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024b; Wang et al., 2024c, a; Yue et al., 2024; Liu et al., 2024b; Xiao et al., 2024). Our work aligns with these self-rewarding VLMs but differs in two key ways: (1) our proposed Image-DPO generates multiple images for a single question-image pair (rather than multiple answers); (2) rather than relying solely on existing images, Image-DPO pipeline creates diverse new images using pre-trained generative models (SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), InstructPix2Pix (Brooks et al., 2023a)). Furthermore, Image-DPO deliberately corrupts images to produce multiple degraded versions that serve as rejected data in DPO training. Concurrent works (Wang et al., 2024a; Xie et al., 2024) explore similar methodologies but lack a benchmark to verify enhanced visual focus, fail to establish theoretical connections between their proposed objective and DPO, and utilize limited image transformations (e.g., only randomly cropping). In contrast, we introduce ViLP (Section 3) to assess visual reasoning and provide theoretical foundations (Appendix B), alongside multi-category image corruptions (semantic editing, Gaussian blurring, pixelation).
@@ 3.1 Design Principles
“What’s the tall animal with the longest neck shown?” Humans readily guess “giraffe” based on learned priors, , yet as shown in the bottom-right of Figure 1, it could be an elephant or camel – where visual reasoning corrects the answer. This highlights a potential shortfall in Vision-Language Models (VLMs), which may over-rely on learned visual language priors instead of true visual reasoning, particularly since VLMs are typically fine-tuned on limited image-text data, which is several orders of magnitude smaller than the trained text corpus (Liu et al., 2024a; Tong et al., 2024). Specifically for the scope of visual language priors in this paper, we target (1) strong language priors that lead VLMs to derive answers solely from text, and (2) potential visual priors causing models to overlook critical uncommon visual cues (e.g., unusual zebra spots in Figure 1).
To evaluate how VLMs handle learned visual language priors, we introduce ViLP, a specialized benchmark of out-of-distribution Question-Image-Answer (QIA) triplets guided by two core principles. First, Text-Only Inference ensures that each question can be answered with high confidence using textual clues alone. Second, Visual Influence requires that the correct answer—sometimes contradicting common sense—only emerges once an out-of-distribution image is considered. By forcing models to integrate both textual and visual information, ViLP reveals whether they truly engage in visual reasoning or merely rely on memorized patterns.
Mathematically, let $Q$ be a question, $I$ an image, and $A={a_{\text{prior}},\dots,a_{\text{test}},\dots}$ the set of possible answers. We define $P(a\mid Q)$ as the probability of answer $a$ given $Q$ alone and $P(a\mid Q,I)$ as the probability given both $Q$ and $I$. We consider a prior model $p$, which may represent either human cognition ($P_{\text{human}}$) or a VLM/LLM’s learned visual-language prior ($P_{\theta}$). For constructing our benchmark, we used the following guidances:
Criterion One: The question $Q$ alone should strongly favor $a_{\text{prior}}$, where $\delta_{1}$ is a high-confidence threshold. $a_{\text{prior}}$ usually satisfies common knowledge, such as “soccer ball is a sphere" and “Einstein created $E=mc^{2}$" (Figure 1).
@@
Criterion Two: With the image $I$, the correct answer shifts to $a_{\text{test}}$, where $\delta_{2}$ is another high-confidence threshold. Also, $a_{\text{test}}$ is significantly different from $a_{\text{prior}}$ to ensure that the image has a substantial impact on the answer, where $D$ is a divergence measure, and $\delta_{3}$ is a threshold indicating significant difference. For instance, the image in the 1st row and the 3rd column of Figure 1 turns the answer to torus.
$P(a_{\text{test}}\mid Q,I)\geq\delta_{2},\,D\big{(}P(a_{\text{prior}}\mid Q),P(a_{\text{test}}\mid Q,I)\big{)}\geq\delta_{3}$ (2)
Criterion Three: The answer $a_{\text{test}}$ should be rare and unlikely from $Q$ alone, while $a_{\text{prior}}$ becomes clearly incorrect when considering $I$. This is enforced by a low-confidence threshold $\delta_{4}$ (e.g., Newton as $a_{\text{test}}$ inferred from the image, contradicting Einstein, shown by the image from the 3rd row and 2nd column of Figure 1).
$P(a_{\text{test}}\mid Q)\leq\delta_{4},\,P(a_{\text{prior}}\mid Q,I)\leq\delta_{4}$ (3)
In constructing ViLP, we use human cognition prior $P_{\text{human}}$ as guidance for its design. We evaluated the behavior of the learned priors of VLMs/LLMs by benchmarking $P_{\theta}$ against the human prior $P_{\text{human}}$.